# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

### v1

@@ Abstract
We investigate a new diffuse-interface model that describes creeping two-phase flows (i.e., flows exhibiting a low Reynolds number), especially flows that permeate a porous medium. The system of equations consists of a Brinkman equation for the volume averaged velocity field as well as a convective Cahn–Hilliard equation with dynamic boundary conditions for the phase-field, which describes the location of the two fluids within the domain. The dynamic boundary conditions are incorporated to model the interaction of the fluids with the wall of the container more precisely. In particular, they allow for a dynamic evolution of the contact angle between the interface separating the fluids and the boundary, and also for a convection-induced motion of the corresponding contact line. For our model, we first prove the existence of global-in-time weak solutions in the case where regular potentials are used in the Cahn–Hilliard subsystem. In this case, we can further show the uniqueness of the weak solution under suitable additional assumptions. Moreover, we further prove the existence of weak solutions in the case of singular potentials. Therefore, we regularize such singular potentials by a Yosida approximation, such that the results for regular potentials can be applied, and eventually pass to the limit in this approximation scheme.
@@ 1   Introduction
The mathematical study of two-phase flows is an important topic in many areas of applied science such as engineering, chemistry and biology. To predict the motion of two fluids, it is crucial to understand how the interface between these fluids evolves. To provide a mathematical description of this interface, two fundamental methods have been developed: the sharp-interface approach and the diffuse-interface approach. In the former, the interface is represented as a hypersurface evolving in the surrounding domain. The occurring quantities (e.g., the velocity fields) are then described by a free boundary problem. In the latter, the fluids are represented by a phase-field function which is expected to attain values close to $1$ in the region where the first fluid is present, and close to $-1$ in the region where the second fluid is located. However, unlike in sharp-interface models, this phase-field does not jump between the values $1$ and $-1$ but exhibits a continuous transition between these values in a thin tubular neighborhood around the boundary between the fluids. This tubular neighborhood is referred to as the diffuse interface and its thickness is proportional to a small parameter $\epsilon>0$. For a comparison of sharp-interface methods and diffuse-interface methods, we refer to [28, 4, 35, 55]. We point out that, even though the sharp-interface and the diffuse-interface approach are conceptually different, they can, in general, be related by the sharp-interface limit in which a parameter related to the thickness of the diffuse interface is sent to zero.
@@
    The no-slip boundary condition on the velocity field makes the model not very suitable to describe general moving contact line phenomena. As the trace of the velocity field at the boundary is fixed to be identically zero, any motion of the contact line of the diffuse interface can be caused only by diffusive but not by convective effects.
@@
associated with the fluids is very small ($\mathrm{Re}\ll 1$). This occurs if the flow speed $u$ and/or the characteristic length $L$ of the flow are small compared to the kinematic viscosity $\nu$. In this situation, it is not necessary to describe the time evolution of the velocity field by the full Navier–Stokes equation. Since advective inertial forces are small compared to viscous forces, the material derivative can be neglected. This leads to the Stokes equation. If a creeping flow through a porous medium is to be considered, an additional term accounting for the permeability needs to be included. The velocity field is then determined by the Brinkman equation.
@@
In system (1), $\Omega\subset{\mathbb{R}}^{d}$ with $d\in\{2,3\}$ is a bounded domain with boundary $\Gamma:=\partial\Omega$, $T>0$ is a given final time, and for brevity, the notation $Q:=\Omega\times(0,T)$ and $\Sigma=\Gamma\times(0,T)$ is used. The vector-valued function ${\boldsymbol{v}}:Q\to{\mathbb{R}}^{d}$ stands for the volume averaged velocity field associated with the fluid mixture and
@@
where $\eta(\varphi)>0$ denotes the dynamic viscosity, and the constants $\varrho>0$ and $\varkappa>0$ stand for the porosity and the intrinsic permeability of the porous medium, respectively. If both $\nu(\varphi)$ and $\lambda(\varphi)$ are positive, (1.1a) is the (quasi-stationary) Brinkman equation which describes the flow through a porous medium. However, if the porosity $\varkappa$ is large compared to the viscosity $\eta(\varphi)$, the function $\lambda(\varphi)$ is very small and can be neglected. In this case we enter the Stokes regime, where no porous media is considered (or the effects of the porous medium are at least negligible). In the formal limit $\varkappa\to\infty$ or $\lambda(\varphi)\to 0$, (1.1a) degenerates to the Stokes equation. In our analysis, we will be able to handle the Brinkman case ($\nu(\varphi)>0$ and $\lambda(\varphi)>0$) and the Stokes case ($\nu(\varphi)>0$ and $\lambda(\varphi)\equiv 0$) simultaneously. On the other hand, if $\lambda(\varphi)$ remains positive and the porosity $\varrho$ is large compared to the dynamic viscosity $\eta(\varphi)$ such that $\nu(\varphi)$ can be neglected, (1.1a) degenerates to Darcy’s law. However, we are not able to handle this case in terms of mathematical analysis as due to the absence of spatial derivatives of the velocity field in (1.1a), we would not obtain enough regularity to define the trace of ${\boldsymbol{v}}$ on the boundary in a reasonable manner.
@@
[…]
@@
In the case $K=0$, the convective bulk-surface Cahn–Hilliard subsystem (1.1c)–(1.1h) is a special case of the one introduced in [40] since for the chemical potential $\mu$, a homogeneous Neumann type boundary condition is imposed in (1.1h). This corresponds to the choice $L=\infty$ in [40]. Therefore, by system (1), we describe a situation where no transfer of material between bulk and boundary occurs. However, its is important that due to the boundary conditions (1.1e)–(1.1i), the model (1) allows for dynamic changes of the contact angle as well as for a convection-induced motion of the contact line. This means that the limitations (L1) and (L2) explained above can be overcome. It is worth mentioning that this setup of dynamic boundary conditions for the Cahn–Hilliard equation (without coupling to a velocity equation) was originally derived in [50] by the Energetic Variational Approach. This system was further investigated in [19, 33, 53, 46]. For similar works on the Cahn–Hilliard equation with Cahn–Hilliard type dynamic boundary conditions, we refer to [49, 47, 34, 42, 14, 17, 18, 64].
In contrast to the model introduced in [40], the phase-fields $\varphi$ and $\psi$ are not just coupled by the trace relation $\varphi|_{\Sigma}=\psi$ on $\Sigma$, but by the more general Robin type coupling condition $K\partial_{\bf n}\varphi=\psi-\varphi$ with $K\geq 0$ (see (1.1g)). This also includes the trace relation via the choice $K=0$. The coupling condition (1.1g) was first used in [16] for an Allen–Cahn type dynamic boundary condition, and later in [46] for a Cahn–Hilliard type dynamic boundary condition. In particular, it was rigorously shown in [46] that the Dirichlet type coupling condition $\varphi|_{\Sigma}=\psi$ on $\Sigma$ can be recovered in the asymptotic limit $K\to 0$. From a physical point of view, the boundary condition (1.1g) with $K>0$ makes sense if the materials on the boundary may be different from those in the bulk. For instance, this might be the case if the materials on the boundary are transformed by chemical reactions. Apart from this, the boundary condition (1.1g) with $K>0$ has a crucial advantage for the mathematical analysis concerning the construction of weak solutions. The reason is that in the case $K=0$, the Dirichlet type coupling condition for the phase-fields already fixes one degree of freedom and is therefore a bad match for the no-mass-flux boundary condition (1.1h)1italic-($_{1}$. This seems to make the direct construction of weak solutions by a Faedo–Galerkin scheme impossible. However, employing (1.1g) with $K>0$, such problems do not arise. Hence, our strategy is to first prove the existence of weak solutions for $K>0$ by a Faedo–Galerkin approach. Afterwards, we construct a weak solution associated with $K=0$ by passing to the limit $K\to 0$ in a suitable sense.
An important property of the system (1) (for any $K\geq 0$) is its thermodynamic consistency with respect to the free energy functional
@@
where $\sigma(K)=K^{-1}$ if $K>0$ and $\sigma(K)=0$ if $K=0$. This means that sufficiently regular solutions of (1) satisfy the energy dissipation law
[…]
@@
The structure of the paper is as follows. In Section 2 we collect some notation, assumptions, preliminaries and important tools as well as the main results of our work. In the case of regular potentials, the Cahn–Hilliard–Brinkman system is analyzed in Section 3. There, we first establish the existence of a weak solution in the case $K>0$, where phase-fields are coupled by a Robin type boundary condition. By sending $K\to 0$, we then obtain the existence of a weak solution also in the case $K=0$, where the phase-fields are coupled by a trace relation. Afterwards, we prove uniqueness in all cases $K\geq 0$ under suitable additional assumptions. Eventually, in Section 4, we show the existence of a weak solution to the Cahn–Hilliard–Brinkman system with singular potentials. This is done based on the results established in Section 3 by approximating the singular potentials by means of Yosida regularizations and finally passing to the limit.
@@ 2.1   Notation
Throughout the manuscript, $\Omega$ is a bounded domain in ${\mathbb{R}}^{d}$, $d\in\{2,3\}$, with Lipschitz boundary $\Gamma:=\partial\Omega$ and ${\bf n}$ is the associated outward unit normal vector field. We write $|\Omega|$ and $|\Gamma|$ to denote the Lebesgue measure of $\Omega$ and the Hausdorff measure of $\Gamma$, respectively. For any given Banach space $X$, we denote its norm by $\left\|\,\cdot\,\right\|_{X}$, its dual space by $X^{*}$ and the duality pairing between $X^{*}$ and $X$ by $\mathopen{\langle}\cdot,\cdot\mathclose{\rangle}_{X}$. Besides, if $X$ is a Hilbert space, we write $(\cdot,\cdot)_{X}$ to denote the corresponding inner product. For every $1\leq p\leq\infty$, $k\geq 0$ and $s>0$, the standard Lebesgue spaces, Sobolev–Slobodeckij spaces and Sobolev spaces defined on $\Omega$ are denoted by $L^{p}(\Omega)$, $W^{k,p}(\Omega)$ and $H^{s}(\Omega)$, and their standard norms are denoted by $\left\|\,\cdot\,\right\|_{L^{p}(\Omega)}$ $\left\|\,\cdot\,\right\|_{W^{k,p}(\Omega)}$ and $\left\|\,\cdot\,\right\|_{H^{s}(\Omega)}$, respectively. It is well known that the spaces $H^{0}(\Omega)=L^{2}(\Omega)$ and $H^{k}(\Omega)=W^{k,2}(\Omega)$ for all $k\in{\mathbb{N}}$ can be identified, and these spaces are Hilbert spaces. The Lebesgue spaces, Sobolev–Slobodeckij spaces and Sobolev spaces on the boundary $\Gamma$ are defined analogously. For more details, we refer to [62, 63]. Moreover, for any Banach spaces $X$ and $Y$, their intersection $X\cap Y$ is also a Banach space subject to the norm
@@
As some spaces will appear very frequently, we introduce the shortcuts:
$H$ $:=L^{2}(\Omega),\quad$ $H_{\Gamma}$ $:=L^{2}(\Gamma),\quad$ $V$ $:=H^{1}(\Omega),\quad$ ${V}_{\Gamma}$ $:=H^{1}(\Gamma),$
${\bf{H}}$ $:=L^{2}(\Omega;{\mathbb{R}}^{d}),\quad$ $\mathbf{H}_{\Gamma}$ $:=L^{2}(\Gamma;{\mathbb{R}}^{d}),\quad$ ${\bf{V}}$ $:=H^{1}(\Omega;{\mathbb{R}}^{d}),$
${\mathbb{H}}$ $:=L^{2}(\Omega;{\mathbb{R}}^{d\times d}).$
We further introduce the spaces of solenoidal (divergence-free) velocity fields:
${\bf H}_{\sigma,{\bf n}}$ :=$:=\big{\{}{\boldsymbol{w}}\in{\bf{H}}:\mathop{\rm div}\nolimits({\boldsymbol{w}})=0\;\text{in $\Omega$},\;{\boldsymbol{w}}|_{\Gamma}\cdot{\bf n}=0\;\text{on $\Gamma$}\big{\}},$
${\bf V}_{\sigma,{\bf n}}$ $:={\bf{V}}\cap{\bf H}_{\sigma,{\bf n}}.$
@@
${\cal V}_{K}$ :=$:=\begin{cases}{\cal V}&\text{if $K>0$},\\ \{(w,w_{\Gamma})\in{\cal V}:w_{\Gamma}={w}|_{\Gamma}\,\,\text{on $\Gamma$}\}&\text{if $K=0$},\end{cases}$
@@
so that ${\cal H}$, ${\cal V}$ and ${\cal V}_{K}$ are Hilbert spaces. For any $v\in{V^{*}}$ and $v_{\Gamma}\in{V}_{\Gamma}^{*}$, we define the generalized mean values by
@@
where $1$ represents the constant function assuming value $1$ in $\Omega$ and on $\Gamma$, respectively. To introduce a weak formulation of (1), it will be useful to define the function

### v2

@@ Abstract
We investigate a new diffuse-interface model that describes creeping two-phase flows (i.e., flows exhibiting a low Reynolds number), especially flows that permeate a porous medium. The system of equations consists of a Brinkman equation for the volume averaged velocity field and a convective Cahn–Hilliard equation with dynamic boundary conditions for the phase-field, which describes the location of the two fluids within the domain. The dynamic boundary conditions are incorporated to model the interaction of the fluids with the wall of the container more precisely. In particular, they allow for a dynamic evolution of the contact angle between the interface separating the fluids and the boundary, and for a convection-induced motion of the corresponding contact line. For our model, we first prove the existence of global-in-time weak solutions in the case where regular potentials are used in the Cahn–Hilliard subsystem. In this case, we can further show the uniqueness of the weak solution under suitable additional assumptions. We further prove the existence of weak solutions in the case of singular potentials. Therefore, we regularize such singular potentials by a Moreau–Yosida approximation, such that the results for regular potentials can be applied, and eventually pass to the limit in this approximation scheme.
@@ 1  Introduction
The mathematical study of two-phase flows is an important topic in many areas of applied science such as engineering, chemistry and biology. To predict the motion of two fluids, it is crucial to understand how the interface between these fluids evolves. To provide a mathematical description of this interface, two fundamental methods have been developed: the sharp-interface approach and the diffuse-interface approach. In the former, the interface is represented as a hypersurface evolving in the surrounding domain. The occurring quantities (e.g., the velocity fields) are then described by a free boundary problem. In the latter, the fluids are represented by a phase-field function which is expected to attain values close to $1$ in the region where the first fluid is present, and close to 1-1 in the region where the second fluid is located. However, unlike in sharp-interface models, this phase-field does not jump between the values $1$ and 1-1 but exhibits a continuous transition between these values in a thin tubular neighborhood around the boundary between the fluids. This tubular neighborhood is referred to as the diffuse interface and its thickness is proportional to a small parameter $\epsilon>0$. For a comparison of sharp-interface methods and diffuse-interface methods, we refer to [34, 6, 45, 70]. We point out that, even though the sharp-interface and the diffuse-interface approach are conceptually different, they can, in general, be related by the sharp-interface limit in which a parameter related to the thickness of the diffuse interface is sent to zero.
@@
    The no-slip boundary condition on the velocity field makes the model not very suitable to describe general moving contact line phenomena. As the trace of the velocity field at the boundary is fixed to be identically zero, any motion of the contact line of the diffuse interface can be caused only by diffusive but not directly by convective effects.
@@
associated with the fluids is very small ($\mathrm{Re}\ll 1$). This occurs if the flow speed $u$ and/or the characteristic length $L$ of the flow are small compared to the kinematic viscosity $\nu$. In this situation, it is not necessary to describe the time evolution of the velocity field by the full Navier–Stokes equation. Since advective inertial forces are small compared to viscous forces, the material derivative can be neglected. This leads to the Stokes equation. If a creeping flow through a porous medium is to be considered, an additional term accounting for the permeability needs to be included. The velocity field is then usually determined by Darcy’s law or the Brinkman equation. For a derivation of these velocity equations via homogenization techniques, we refer, for instance, to [10, 57, 66, 53, 37] and the references therein.
@@
In system (1.1), $\Omega\subset{\mathbb{R}}^{d}$ with $d\in\{2,3\}$ is a bounded domain with boundary $\Gamma:=\partial\Omega$, $T>0$ is a given final time, and for brevity, the notation $Q:=\Omega\times(0,T)$ and $\Sigma=\Gamma\times(0,T)$ is used. The vector-valued function ${\boldsymbol{v}}:Q\to{\mathbb{R}}^{d}$ stands for the volume averaged velocity field associated with the fluid mixture and
@@
where $\eta(\varphi)>0$ denotes the dynamic viscosity, and the constants $\varrho>0$ and $\varkappa>0$ stand for the porosity and the intrinsic permeability of the porous medium, respectively. If both $\nu(\varphi)$ and $\lambda(\varphi)$ are positive, (1.1a) is the (quasi-stationary) Brinkman equation which describes the flow through a porous medium. It is equipped with the incompressibility condition (1.1b) and the inhomogeneous Navier-slip boundary condition (1.1i) in which the slip parameter $\gamma(\psi)\geq 0$ may depend on the surface phase-field $\psi$. However, if the porosity $\varkappa$ is large compared to the viscosity $\eta(\varphi)$, the function $\lambda(\varphi)$ is very small and can be neglected. In this case we enter the Stokes regime, where no porous media is considered (or the effects of the porous medium are at least negligible). In the formal limit $\varkappa\to\infty$ or $\lambda(\varphi)\to 0$, (1.1a) degenerates to the Stokes equation. In our analysis, we will be able to handle the Brinkman case ($\nu(\varphi)>0$ and $\lambda(\varphi)>0$) and the Stokes case ($\nu(\varphi)>0$ and $\lambda(\varphi)\equiv 0$) simultaneously. On the other hand, if $\lambda(\varphi)$ remains positive and the porosity $\varrho$ is large compared to the dynamic viscosity $\eta(\varphi)$ such that $\nu(\varphi)$ can be neglected, (1.1a) degenerates to Darcy’s law. However, we are not able to handle this case in terms of mathematical analysis as due to the absence of spatial derivatives of the velocity field in (1.1a), we would not obtain enough regularity to define the trace of $\boldsymbol{v}$} on the boundary in a reasonable manner.
@@
for all $s\in(-1,1)$. Here, $\Theta>0$} is the absolute temperature of the mixture, and $\Theta_{c}$ is a critical temperature such that phase separation will occur in case $0<\Theta<\Theta_{c}$. The logarithmic potential is classified as a singular potential since its derivative $F^{\prime}$ diverges to $\pm\infty$ when its argument approaches $\pm 1$. It is often approximated by a polynomial double-well potential
@@
In the case $K=0$, the convective bulk-surface Cahn–Hilliard subsystem (1.1c)–(1.1h) is a special case of the one introduced in [51] since for the chemical potential $\mu$, a homogeneous Neumann type boundary condition is imposed in (1.1h). This corresponds to the choice $L=\infty$ in [51]. Therefore, by system (1.1), we describe a situation where no transfer of material between bulk and boundary occurs. However, its is important that due to the boundary conditions (1.1e)–(1.1i), the model (1.1) allows for dynamic changes of the contact angle as well as for a convection-induced motion of the contact line. This means that the limitations (L1) and (L2) explained above can be overcome. It is worth mentioning that this setup of dynamic boundary conditions for the Cahn–Hilliard equation (without coupling to a velocity equation) was originally derived in [64] by the Energetic Variational Approach. This system was further investigated in [24, 43, 67, 59]. For similar works on the Cahn–Hilliard equation with Cahn–Hilliard type dynamic boundary conditions, we refer to [62, 60, 44, 54, 19, 22, 23, 79].
In contrast to the model introduced in [51], the phase-fields $\varphi$ and $\psi$ are not just coupled by the trace relation $\varphi|_{\Sigma}=\psi$ on $\Sigma$, but by the more general Robin type coupling condition $K\partial_{\bf n}\varphi=\psi-\varphi$ with $K\geq 0$ (see (1.1g)). This also includes the trace relation via the choice $K=0$. The coupling condition (1.1g) was first used in [21] for an Allen–Cahn type dynamic boundary condition, and later in [59] for a Cahn–Hilliard type dynamic boundary condition. In particular, it was rigorously shown in [59] that the Dirichlet type coupling condition $\varphi|_{\Sigma}=\psi$ on $\Sigma$ can be recovered in the asymptotic limit $K\to 0$. From a physical point of view, the boundary condition (1.1g) with $K>0$ makes sense if the materials on the boundary may be different from those in the bulk. For instance, this might be the case if the materials on the boundary are transformed by chemical reactions.
An important property of the system (1.1) (for any $K\geq 0$) is its thermodynamic consistency with respect to the free energy functional
@@
where $\sigma(K)=K^{-1}$ if $K>0$ and $\sigma(K)=0$ if $K=0$. This means that sufficiently regular solutions of (1.1) satisfy the energy dissipation law
$\begin{aligned} \frac{\mathrm{d}}{\mathrm{d}t}E_{K}(\varphi,\psi)&=-\int_{\Omega}\lambda(\varphi)|{\boldsymbol{v}}|^{2}-\int_{\Omega}M_{\Omega}(\varphi)|\nabla\mu|^{2}-\int_{\Gamma}M_{\Gamma}(\psi)|\nabla_{\Gamma}\theta|^{2}\\ &\quad-2\int_{\Omega}\nu(\varphi)|\mathbf{D}{\boldsymbol{v}}|^{2}-\int_{\Gamma}\gamma(\psi)|{\boldsymbol{v}}|^{2},\end{aligned}$
@@
Goals and novelties of this paper.
After collecting some notations, assumptions, preliminaries and important tools, we introduce our main results in Section 2. Let us now briefly discuss the results with emphasis on the objectives of this work.
  1. (I)
    Our first main goal is to establish the weak well-posedness to system (1.1) in the case of regular potentials $F$ and $G$ for all choices $K\geq 0$.
[…]
    In previous works in the literature (e.g., [43, 59, 60]), where the bulk-surface Cahn–Hilliard system was considered without any velocity field, such issues of unmatched boundary conditions could be overcome by exploiting the gradient flow structure of the system and employing a minimizing movement scheme for the construction of a weak solution. However, as in our case the bulk-surface Cahn–Hilliard system is coupled to a velocity equation, the whole system is not a gradient flow anymore and therefore, a minimizing movement technique is not applicable.
[…]
    The existence of a weak solution to (1.1) for any $K>0$ is stated in Theorem 2.6. For the proof, we use a semi-Galerkin scheme, where only the phase-fields and the chemical potentials are discretized using a Galerkin scheme, but the corresponding velocity field is obtained by directly solving the Brinkman subsystem on the continuous level. For a Cahn–Hilliard–Brinkman model without dynamic boundary conditions, such an approach had already been employed in [35]. Based on this ansatz, a sequence of approximate solutions can be constructed by means of the Cauchy–Peano theorem and eventually, after deriving suitable uniform bounds, we show that this sequence converges to a weak solution of (1.1) with $K>0$. A posteriori, we establish higher regularity of the phase-field functions provided that the domain is sufficiently smooth. The corresponding proof can be found in Subsection 3.1.
    In Theorem 2.7, we construct a weak solution to (1.1) in the case $K=0$ by using the result of Theorem 2.6, and passing to the limit $K\to 0$ on the level of weak solutions. A posteriori, as in the case $K>0$, higher regularity of the phase-fields can be shown if the domain is sufficiently regular. The proof is carried out in Subsection 3.2.
  2. (II)
    The second main goal is to prove the uniqueness as well as the continuous dependence on the initial data of the weak solution to (1.1) with regular potentials $F$ and $G$ in all cases $K\geq 0$, provided that the mobility functions and the viscosity function are constant and the domain is sufficiently regular.
    The result is stated in Theorem 2.8 and the proof is presented in Subsection 3.3. To prove the assertion, we adapt the uniqueness proof devised in [51] for a Navier–Stokes–Cahn–Hilliard model to our situation. We point out that the proof in [51] only worked in two dimensions, whereas our new proof for the system (1.1) is also valid in three dimensions. Moreover, our new proof requires a certain bulk-surface chain rule for time derivatives that is established in Proposition A.1 in the Appendix.
  3. (III)
    Our third main result is to prove the existence of a weak solution to system (1.1) in the case of possibly singular potentials $F$ and $G$ for all choices $K\geq 0$.
    The result is stated in Theorem 2.10 and the proof is presented in Section 4. The idea of the proof is to approximate the convex parts of the singular potentials $F$ and $G$ by regular potentials $F_{\varepsilon}$ and $G_{\varepsilon}$ (with $\varepsilon>0$) that are constructed via a Moreau–Yosida regularization. For the approximate regular potentials, the existence of a corresponding weak solutions is already known from Theorem 2.6. The most technical part of the proof of Theorem 2.10 is to derive uniform bounds on the potential terms involving $F_{\varepsilon}^{\prime}$ and $G_{\varepsilon}^{\prime}$. This requires a certain condition on the singular potentials, namely that $G^{\prime}$ dominates $F^{\prime}$ in a suitable way (cf (S2)). Eventually, by passing to the limit $\varepsilon\to 0$, the claim of Theorem 2.10 is established.
@@ 2.1  Notation
Throughout the manuscript, $\Omega$ is a bounded domain in ${\mathbb{R}}^{d}$, $d\in\{2,3\}$, with Lipschitz boundary $\Gamma:=\partial\Omega$ and $\bf n$} is the associated outward unit normal vector field. We write $|\Omega|$ and $|\Gamma|$ to denote the Lebesgue measure of $\Omega$ and the Hausdorff measure of $\Gamma$, respectively. For any given Banach space $X$, we denote its norm by $\left\|\,\cdot\,\right\|_{X}$, its dual space by $X^{*}$ and the duality pairing between $X^{*}$ and $X$ by $\mathopen{\langle}\cdot,\cdot\mathclose{\rangle}_{X}$. Besides, if $X$ is a Hilbert space, we write $(\cdot,\cdot)_{X}$ to denote the corresponding inner product. For every $1\leq p\leq\infty$, $k\geq 0$ and $s>0$, the standard Lebesgue spaces, Sobolev–Slobodeckij spaces and Sobolev spaces defined on $\Omega$ are denoted by $L^{p}(\Omega)$, $W^{k,p}(\Omega)$ and $H^{s}(\Omega)$, and their standard norms are denoted by $\left\|\,\cdot\,\right\|_{L^{p}(\Omega)}$ $\left\|\,\cdot\,\right\|_{W^{k,p}(\Omega)}$ and $\left\|\,\cdot\,\right\|_{H^{s}(\Omega)}$, respectively. It is well known that the spaces $H^{0}(\Omega)=L^{2}(\Omega)$ and $H^{k}(\Omega)=W^{k,2}(\Omega)$ for all $k\in{\mathbb{N}}$ can be identified, and these spaces are Hilbert spaces. The Lebesgue spaces, Sobolev–Slobodeckij spaces and Sobolev spaces on the boundary $\Gamma$ are defined analogously. For more details, we refer to [77, 78]. We usually utilize bold letters to represent spaces for vector- or matrix-valued functions. For example, we denote ${\bf{L}}^{p}(\Omega)$ instead of writing $L^{p}(\Omega;{\mathbb{R}}^{d})$ or $L^{p}(\Omega;{\mathbb{R}}^{d\times d})$, and so on. Moreover, for any Banach spaces $X$ and $Y$, their intersection $X\cap Y$ is also a Banach space subject to the norm
@@
As some spaces will appear very frequently, we introduce the following shortcuts:
$\begin{aligned} &H:=L^{2}(\Omega),\quad&&H_{\Gamma}:=L^{2}(\Gamma),\quad&&V:=H^{1}(\Omega),\quad&&{V}_{\Gamma}:=H^{1}(\Gamma),\\ &{\bf{H}}:=L^{2}(\Omega;{\mathbb{R}}^{d}),\quad&&\mathbf{H}_{\Gamma}:=L^{2}(\Gamma;{\mathbb{R}}^{d}),\quad&&{\bf{V}}:=H^{1}(\Omega;{\mathbb{R}}^{d}),\quad&&{\mathbb{H}}:=L^{2}(\Omega;{\mathbb{R}}^{d\times d}),\end{aligned}$
${\bf H}_{\sigma,{\bf n}}:=\big{\{}{\boldsymbol{w}}\in{\bf{H}}:\mathop{\rm div}\nolimits({\boldsymbol{w}})=0\;\text{in $\Omega$},\;{\boldsymbol{w}}|_{\Gamma}\cdot{\bf n}=0\;\text{on $\Gamma$}\big{\}},\quad{\bf V}_{\sigma,{\bf n}}:={\bf{V}}\cap{\bf H}_{\sigma,{\bf n}},$
${\bf{V}}_{\bf n}:=\big{\{}{\boldsymbol{w}}\in{\bf{V}}:{\boldsymbol{w}}|_{\Gamma}\cdot{\bf n}=0\;\text{on $\Gamma$}\big{\}},\quad W_{\bf n}:=\big{\{}w\in H^{2}(\Omega):\partial_{\bf n}w=0\;\text{on $\Gamma$}\big{\}}.$
@@
${\cal V}_{K}$ $:=\begin{cases}{\cal V}&\text{if $K>0$},\\ \{(w,w_{\Gamma})\in{\cal V}:w_{\Gamma}={w}|_{\Gamma}\,\,\text{on $\Gamma$}\}&\text{if $K=0$},\end{cases}$
@@
[…]
${\cal V}^{*}=V^{*}\times V_{\Gamma}^{*},$
${\cal V}_{K}\subset{\cal V}\quad\text{and}\quad{\cal V}^{*}\subset{\cal V}_{K}^{*}\quad\text{if $K=0$},$
${\cal V}_{K}={\cal V}\quad\text{and}\quad{\cal V}^{*}={\cal V}_{K}^{*}\quad\text{if $K>0$}.$
For any $v\in{V^{*}}$ and $v_{\Gamma}\in{V}_{\Gamma}^{*}$, we define the generalized mean values by
@@
where $1$ represents the constant function assuming value $1$ in $\Omega$ and on $\Gamma$, respectively. To introduce a weak formulation of (1.1), it will be useful to define the function